A movie theater shows films for a community of 10000 people shows films during weekends. Right now, the price per ticket is $17.50. In the past, when they increased or decreased the price per ticket, they discovered that for every dollar (or fraction) that the price was increased or decreased, the attendance decreased or increased proportionally by 200 people. The theater owner pays the Film Distribution Company $10 (incremental cost) in royalties per person who views the film. Can be Hand written or in excel just need to show work. Find

  1. The willingness-to-pay distribution of the movie goers. 
  2. The average price they would be willing to pay.
  3. The price per ticket that maximizes revenue. 
  4. Should the theater owner increase or decrease the price per ticket if she wants to maximize contribution?
  5. Find the consumer surplus if the tickets are priced at the price that maximizes contribution.
1
Expert's answer
2021-03-11T07:20:46-0500

"solution"

(A)

Willing to pay(WTP), sometimes shortened as WTP is the maximum value a customer is

WTP for goods or services.


WTP="200\times10000=2,000,000" pay distribution of movie goers


B) The price which a customer pays for a the line is always less than what he is willing to pay for it, so that the satisfaction which he gets from its investment is more than the price paid for it and thus he begins a surplus content which Marshall calls Consumer's Surplus.

the average price they are willing to pay hence will be "\$17.50"


C) Let "x" be the price of the ticket.

"If\ x=10, 2000" seats last sold and that 100 fewer seats are sold for each dollar rise in

price.

"seats\ sold =2000-100(x+10)"

If we add this to the cost of the ticket, we

can find the total revenue RR:

"R=x(2000-100(x-10))\\

R=2000x-100x(x-10)\\

R=2000x-100x2+1000x\\

R=3000x-100x2\\"

Because we can run this office for pay for any monotonically rising analytic use and their maximum will occur at the same value of "x."

"R=30x-x2"

Taking the derivative concerning"\ x" , we

have:

"R=30-2x"

and setting this to zero, we have:

"0=30-2x"

"x=15"

Finally, we can verify that this is a maximum and not a minimum because the second the derivative of our revenue function is negative. Therefore, you'll want to price your tickets at

"x=15" .


D) Yes, the business owner should improve or lower the cost per ticket if she needs to maximize donation


E) There is a business method that is used to the consumer surplus is the separation of the highest users would pay including the actual price they pay.


"\$17.50-\$10=\$7.5" is the consumer surplus